# Environmental Breakdown of Topological Interface States in Armchair Graphene Nanoribbon Heterostructures: Is it true?

### v1

@@ Abstract
We investigate the environmental stability of topological interface states (IFs) in two sandwich nanostructures, BNNR/AGNRH/BNNR and BNNR/AGNRH/NBNR, where AGNRH denotes an armchair graphene nanoribbon heterostructure and BNNR (NBNR) represents a boron nitride nanoribbon. The former corresponds to a same-topology configuration, whereas the latter realizes a reverse-topology configuration. Using a bulk boundary perturbation approach, we show that in BNNR/AGNRH/BNNR the IFs are destroyed by chirality breaking induced by symmetric BN environments at both interfaces. In contrast, the IFs in the reverse-topology structure remain robust against lateral interface interactions from BN atoms. Transport calculations further demonstrate that the surviving IFs in BNNR/AGNRH/NBNR exhibit the characteristic behavior of topological double quantum dots, with an enhanced interdot hopping strength compared with vacuum boundary conditions. These results reveal that BN environments can either suppress or reinforce topological interface states, depending critically on the topology of the surrounding nanoribbons.
Introduction–Bottom-up synthesis techniques enable the fabrication of graphene nanoribbons (GNRs) with atomic precision [Cai, 1DJRizzo, 10]. Compared with zigzag GNRs (ZGNRs), armchair GNRs (AGNRs) exhibit tunable electronic phases [LlinasJP, 5], making them promising candidates for nanoelectronic applications [WangHM, 9]. More complex GNR architectures, including 9-7-9 and 7-9-7 armchair graphene nanoribbon heterostructures (AGNRHs) [DJRizzo, 10] and Janus GNR segments [SongST, 11], have also been experimentally realized. Scanning tunneling microscopy measurements have confirmed the presence of topological states (TSs) in these systems, such as end states (ESs) and interface states (IFs) in 9-7-9 and 7-9-7 AGNRHs [DJRizzo, 10].
@@
Surprisingly, we find that BNNR/AGNRH/NBNR structures with reverse topology preserve the functionality of topological double quantum dots (TDQDs), in which a single energy level associated with each quantum dot remains well isolated from the bulk states. Moreover, the electron hopping strength between the TDQDs is significantly enhanced compared with the vacuum boundary condition. This behavior is highly unusual when contrasted with conventional semiconductor double quantum dot (DQD) systems [Ono, 19,vanderWiel, 20], whose operation is typically restricted to ultra-low temperatures due to the small energy separation between bound and continuum states. Nanoelectronic devices based on TDQDs therefore offer a promising route toward robust high-temperature quantum functionality.
Figure 1: Schematic illustration of an armchair graphene nanoribbon heterostructure (AGNRH) laterally embedded in a hexagonal boron nitride (BN) sheet, forming two distinct configurations: (a) a same-topology structure, 2-BNNR/9-7-9-AGNRH/2-BNNR, and (b) a reverse-topology structure, 2-BNNR/9-7-9-AGNRH/2-NBNR. White and blue sites denote boron and nitrogen atoms, respectively. $R_{1}$ and $R_{2}$ represent the unit cells (u.c.) of the 9-AGNR/2-BN and 7-AGNR/3-BN junctions, respectively. The tunable parameter $t_{in}$ characterizes the interfacial coupling between the AGNRH and the BNNRs. The symbols $\Gamma_{L}$ ($\Gamma_{R}$) represent the electron tunneling rates between the left (right) electrode and the leftmost (rightmost) atoms at the zigzag edges.
Calculation Methodology–To investigate the transport properties of the BNNR/AGNRH/NBNR heterojunction connected to metallic electrodes, we employ a combination of a tight-binding model and the Green’s function technique [QiaoZH, 17]. The system Hamiltonian, illustrated in Fig. 1(b), is written as $H=H_{0}+H_{GNR}$, where $H_{0}$ describes the electrodes and their coupling to the nanoribbon system, and $H_{GNR}$ represents the BNNR/9-7-9-AGNRH/NBNR structure.
@@
Here, $E_{\ell,j}$ denotes the on-site energy of the orbital at site $(\ell,j)$, while $d^{\dagger}_{\ell,j}$ and $d_{\ell,j}$ are the electron creation and annihilation operators, respectively. The parameter $t_{(\ell,j),(\ell^{\prime},j^{\prime})}$ characterizes the hopping amplitude between sites $(\ell^{\prime},j^{\prime})$ and $(\ell,j)$. For the BNNR/AGNRH/NBNR system, we assign on-site energies $E_{B}=\Delta_{B}$, $E_{N}=-\Delta_{N}$, and $E_{C}=0$ for boron, nitrogen, and carbon atoms, respectively. Variations in hopping strengths between different atomic species are neglected, and the condition $E_{B}=-E_{N}=\Delta=2.7$ eV is adopted for simplicity [DingY, 21KuoJL, 24]. The nearest-neighbor hopping energy is set to $t_{(\ell,j),(\ell^{\prime},j^{\prime})}=t_{pp\pi}=2.7$ eV.
To quantify interfacial interactions induced by the BN environment, we introduce an inter-nanoribbon hopping parameter $t_{in}$ based on bulk boundary perturbation approach[DavidKuo, 12], as illustrated in Fig. 1(a). When $t_{in}=0$, the BNNR/AGNRH/BNNR (or BNNR/AGNRH/NBNR) structure is decoupled into three isolated segments–the bottom BNNR, the central 9-7-9-AGNRH, and the top BNNR (or NBNR)–each subject to vacuum boundary conditions. Varying $t_{in}$ effectively models different strengths of lateral interface interactions arising from surrounding two-dimensional insulating environments [PerepichkaDF, 25].
@@
Results and discussion–Interface-induced modifications of armchair AGNRs embedded in BN sheets have been experimentally observed and shown to alter the band gaps between conduction and valence subbands [Levendorf, 26GengDC, 29]. To elucidate how BN environments affect the topological IFs of 9-7-9 -AGNRHs, we first examine the influence of BNNRs with different topologies on individual 7-AGNR and 9-AGNR with $R_{1}$ unit cells. Figures 2(a) and 2(b) show the calculated band structures of 4-BNNR/7-AGNR/4-BNNR and 4-BNNR/9-AGNR/4-BNNR, corresponding to same-topology configurations. The resulting band gaps are $1.22$ eV and $1.19$ eV, respectively, compared with $1.0$ eV and $0.9$ eV for isolated 7-AGNR and 9-AGNR under vacuum boundary conditions. These results indicate that BN embedding induces only weak perturbations to the bulk electronic structure when the surrounding BNNRs share the same topology[JungJ, 18,KuoJL, 24].
In contrast, for inverse-topology configurations shown in Figs. 2(c) and 2(d), the band gaps of 4-BNNR/7-AGNR/4-NBNR and 4-BNNR/9-AGNR/4-NBNR are significantly reduced to $0.394$ eV and $0.454$ eV, respectively. This pronounced gap renormalization highlights the strong influence of BN environments when the bottom BNNR has opposite topology to the top NBNR. These tight-binding results are in good agreement with previous DFT calculations [JungJ, 18,KuoJL, 24], and the small residual differences suggest that higher-order hopping processes play a negligible role in the low-energy conduction and valence subbands. We further find that varying the BNNR width ($n=6,8,10,\ldots$) does not qualitatively alter the band-edge dispersions, justifying the use of 4-BNNRs in the following analysis.
@@
In realistic experiments, AGNR and AGNRH segments have finite lengths, typically below $20$ nm[DJRizzo, 10]. We therefore examine the discrete energy spectra of finite 7-AGNR and 9-7-9 AGNRH segments embedded in BN environments with same and inverse topology, as shown in Fig. 3. Figures 3(a) and 3(b) display the energy levels of 4-BNNR/7-AGNR/4-BNNR and 4-BNNR/9-7-9-AGNRH/4-BNNR as functions of the interfacial hopping strength $t_{in}$ for the same-topology case. Owing to the finite size of the 7-AGNR segment $(N=15,M=80)$, two in-gap energy levels labeled $\Sigma_{ES,C}$ and $\Sigma_{ES,V}$ emerge from ESs, as shown in Fig. 3(a). In the absence of BN coupling ($t_{in}=0$), these ESs form near-zero-energy modes. As $t_{in}$ increases, $\Sigma_{ES,C}$ and $\Sigma_{ES,V}$ are strongly shifted away from zero energy, whereas the bulk states remain largely unaffected, consistent with the infinite-length results in Fig. 2.
Figure 3(b) reveals four in-gap energy levels originating from two ESs and two IFs in the 9-7-9 AGNRH segment. The IFs arise from the effective end states of the central 7-AGNR segment with $R_{2}$ unit cells, as discussed in our previous work[DavidKuo, 12]. Within the physically relevant range $0.8~t\leq t_{in}\leq t$, corresponding to realistic BN coupling strengths, the IF levels $\Sigma_{IF,C(V)}$ are no longer well isolated from the bulk states. This behavior contrasts sharply with the $t_{in}=0$ limit and reflects the breakdown of topological protection due to chirality mixing induced by the symmetric BN environment.
The situation changes dramatically in the inverse-topology configuration shown in Figs. 3(c) and 3(d), where finite 7-AGNR and $9_{6}$-$7_{8}$-$9_{6}$ AGNRH segments are embedded between a BNNR and an NBNR. Although the bulk band gap ($E_{g}=E_{C}-E_{V}$) is strongly reduced for $0.8~t\leq t_{in}\leq t$, the energy levels associated with ESs and IFs in the 9-7-9 AGNRH remain nearly unchanged and well separated from the bulk continuum. This robustness is highly nontrivial, demonstrating that inverse topology effectively protects the IFs against strong environmental perturbations.
@@
To clarify the contrasting behaviors observed in Figs. 3(b) and 3(d), we plot the charge densities of $\Sigma_{IF,C}$ for different $t_{in}$ values in Fig. 4. For $t_{in}=0$, shown in Fig. 4(a), the AGNRH is completely decoupled from the BN environment, corresponding to vacuum boundary conditions. The state $\Sigma_{IF,C}$ arises from the coupling between a left IF with A-sublattice chirality ($\Psi_{IF,A}$) and a right IF with B-sublattice chirality ($\Psi_{IF,B}$), resulting in spatially separated charge densities and long-range hopping characteristic of topological states. When a small coupling $t_{in}=0.2~t$ is introduced, the IF wave functions partially penetrate into the BN regions [Fig. 4(b)], leading to weak chirality perturbations. Even for strong coupling $t_{in}=0.8~t$, shown in Fig. 4(c), the IF wave functions remain predominantly sublattice polarized, with only minor overlap between A and B sublattices.
In stark contrast, Figs. 4(d)-4(f) show that for same-topology configurations the IF charge densities become strongly polarized and localized at a single interface, accompanied by substantial overlap between opposite chiralities. This behavior indicates that the symmetric BN environment induces an effective lateral field along the armchair direction, breaking chirality and destroying the topological nature of the IFs.

### v2

@@ Abstract
We theoretically investigate the stability and transport properties of topological interface states (IFs) in 9-7-9 and 15-13-15 armchair graphene nanoribbon heterostructures (AGNRHs) laterally embedded in boron nitride (BN) sheets. Two configurations, $n$-BNNR/AGNRH/$n$-BNNR and $n$-BNNR/AGNRH/$n$-NBNR, corresponding to same-topology and reverse-topology BN environments, are examined within a tight-binding framework. Using a bulk boundary perturbation approach, we show that in BNNR/AGNRH/BNNR the IFs are destroyed by chirality breaking induced by symmetric BN environments at both interfaces. In contrast, the IFs in the reverse-topology structure remain robust against lateral interface interactions from BN atoms. Transport calculations further demonstrate that the surviving IFs in BNNR/AGNRH/NBNR exhibit the characteristic behavior of topological double quantum dots, with an enhanced interdot hopping strength compared with vacuum boundary conditions. These results reveal that BN environments can either suppress or reinforce topological interface states, depending critically on the topology of the surrounding nanoribbons.
I Introduction
Bottom-up synthesis techniques enable the fabrication of graphene nanoribbons (GNRs) with atomic precision [Cai, 1DJRizzo, 10]. Compared with zigzag GNRs (ZGNRs), armchair GNRs (AGNRs) exhibit tunable electronic phases [LlinasJP, 5], making them promising candidates for nanoelectronic applications [WangHM, 9]. More complex GNR architectures, including 9-7-9 and 7-9-7 armchair graphene nanoribbon heterostructures (AGNRHs) [DJRizzo, 10] and Janus GNR segments [SongST, 11], have also been experimentally realized. Scanning tunneling microscopy measurements have confirmed the presence of topological states (TSs) in these systems, such as end states (ESs) and interface states (IFs) in 9-7-9 and 7-9-7 AGNRHs [DJRizzo, 10].
@@ I Introduction
Surprisingly, we find that BNNR/AGNRH/NBNR structures with reverse topology preserve the functionality of topological double quantum dots (TDQDs), in which a single energy level associated with each quantum dot remains well isolated from the bulk states. Moreover, the electron hopping strength between the TDQDs is significantly enhanced compared with the vacuum boundary condition. This behavior is highly unusual when contrasted with conventional semiconductor DQD systems [Ono, 19ZajacDM, 26], whose operation is typically restricted to ultra-low temperatures due to the small energy separation between bound and continuum states. Nanoelectronic devices based on TDQDs therefore offer a promising route toward robust high-temperature quantum functionality.
Figure 1: Schematic illustration of an armchair graphene nanoribbon heterostructure (AGNRH) laterally embedded in a hexagonal boron nitride (BN) sheet, forming two distinct configurations: (a) a same-topology structure, 2-BNNR/9-7-9-AGNRH/2-BNNR, and (b) a reverse-topology structure, 2-BNNR/9-7-9-AGNRH/2-NBNR. White and blue sites denote boron and nitrogen atoms, respectively. $R_{1}$ and $R_{2}$ represent the unit cells (u.c.) of the 2-BNNR/9-AGNR/2-BNNR and 3-BNNR/7-AGNR/3-BNNR junctions, respectively. The tunable parameter $t_{in}$ characterizes the interfacial coupling between the AGNRH and the BNNRs. The symbols $\Gamma_{L}$ ($\Gamma_{R}$) represent the electron tunneling rates between the left (right) electrode and the leftmost (rightmost) atoms at the zigzag edges.
II Calculation Methodology
To investigate the transport properties of the BNNR/AGNRH/NBNR heterojunction connected to metallic electrodes, we employ a combination of a tight-binding model and the Green’s function technique [QiaoZH, 17]. The system Hamiltonian, illustrated in Fig. 1(b), is written as $H=H_{0}+H_{GNR}$, where $H_{0}$ describes the electrodes and their coupling to the nanoribbon system, and $H_{GNR}$ represents the BNNR/AGNRH/NBNR structure.
@@ II Calculation Methodology
Here, $E_{\ell,j}$ denotes the on-site energy of the orbital at site $(\ell,j)$, while $d^{\dagger}_{\ell,j}$ and $d_{\ell,j}$ are the electron creation and annihilation operators, respectively. The parameter $t_{(\ell,j),(\ell^{\prime},j^{\prime})}$ characterizes the hopping amplitude between sites $(\ell^{\prime},j^{\prime})$ and $(\ell,j)$. For the BNNR/AGNRH/NBNR system, we assign on-site energies $E_{B}=\Delta_{B}$, $E_{N}=-\Delta_{N}$, and $E_{C}=0$ for boron, nitrogen, and carbon atoms, respectively. Variations in hopping strengths between different atomic species are neglected, and the condition $E_{B}=-E_{N}=\Delta=2.7$ eV is adopted for simplicity [DingY, 27KuoJL, 30]. The nearest-neighbor hopping energy is set to $t_{(\ell,j),(\ell^{\prime},j^{\prime})}=t_{pp\pi}=2.7$ eV.
To quantify interfacial interactions induced by the BN environment, we introduce an inter-nanoribbon hopping parameter $t_{in}$ based on bulk boundary perturbation approach[DavidKuo, 12], as illustrated in Fig. 1(a). When $t_{in}=0$, the BNNR/AGNRH/BNNR (or BNNR/AGNRH/NBNR) structure is decoupled into three isolated segments–the lower BNNR, the central 9-7-9-AGNRH, and the upper BNNR (or NBNR)–each subject to vacuum boundary conditions. Varying $t_{in}$ effectively models different strengths of lateral interface interactions arising from surrounding two-dimensional insulating environments [PerepichkaDF, 31].
@@
III Results and discussion
Interface-induced modifications of AGNRs embedded in BN sheets have been experimentally observed and shown to alter the band gaps between conduction and valence subbands [Levendorf, 32GengDC, 35]. To elucidate how BN environments affect the topological IFs of 9-7-9 -AGNRHs, we first examine the influence of BNNRs with different topologies on individual 7-AGNR and 9-AGNR with $R_{1}$ unit cells. Figures 2(a) and 2(b) show the calculated band structures of 4-BNNR/7-AGNR/4-BNNR and 4-BNNR/9-AGNR/4-BNNR, corresponding to same-topology configurations. The resulting band gaps are $1.22$ eV and $1.19$ eV, respectively, compared with $1.0$ eV and $0.9$ eV for isolated 7-AGNR and 9-AGNR under vacuum boundary conditions.
For same-topology BN lateral embedding, both the lower and upper BN edges induce sublattice potentials of the same sign. Consequently, the boundary condition imposed by BN resembles that of a vacuum (infinite potential barrier), except that the BN case corresponds to a finite potential barrier. This explains why the band-gap differences between vacuum and BN boundary conditions are relatively small. Indeed, when $\Delta$ is set to a sufficiently large value (e.g., $\Delta=100~\mathrm{eV}$), the two results become nearly indistinguishable. These findings indicate that BN lateral embedding introduces only weak perturbations to the bulk electronic structure when the surrounding BNNRs share the same topology [JungJ, 18,KuoJL, 30].
In contrast, for inverse-topology configurations shown in Figs. 2(c) and 2(d), the band gaps of 4-BNNR/7-AGNR/4-NBNR and 4-BNNR/9-AGNR/4-NBNR are significantly reduced to $0.394$ eV and $0.454$ eV, respectively. This pronounced gap renormalization highlights the strong influence of the BN environment when the lower BNNR has an opposite topology to the upper NBNR. In the inverse-topology case, the sublattice potentials induced at the lower and upper BN edges have opposite signs. Such a boundary condition differs fundamentally from that of AGNRs under vacuum boundary conditions, leading to substantial modifications of the band-edge states.
These tight-binding results are in good agreement with previous DFT calculations [JungJ, 18,KuoJL, 30]. The small residual discrepancies suggest that higher-order hopping processes play a negligible role in determining the low-energy conduction and valence subbands. Furthermore, we find that varying the BNNR width ($n=6,8,10,\ldots$) does not qualitatively modify the band-edge dispersions, thereby justifying the use of 4-BNNRs in the following analysis.
@@ III Results and discussion
In realistic experiments, AGNR and AGNRH segments possess finite lengths, typically below $20$ nm[DJRizzo, 10]. We therefore examine the discrete energy spectra of finite 7-AGNR and 9-7-9 AGNRH segments embedded in BN environments with same and inverse topology, as shown in Fig. 3. Figures 3(a) and 3(b) display the energy levels of 4-BNNR/7-AGNR/4-BNNR and 4-BNNR/9-7-9-AGNRH/4-BNNR as functions of the interfacial hopping strength $t_{in}$ for the same-topology case. Due to the finite size of the 4-BNNR/7-AGNR/4-BNNR segment $(N=15,M=80)$, two in-gap energy levels, denoted as $\Sigma_{ES,C}$ and $\Sigma_{ES,V}$, emerge from the ESs, as shown in Fig. 3(a). In the absence of BN coupling ($t_{in}=0$), these ESs form near-zero-energy modes. As $t_{in}$ increases, $\Sigma_{ES,C}$ and $\Sigma_{ES,V}$ shift markedly away from zero energy, whereas the bulk-derived states remain largely unchanged. This behavior is consistent with the infinite-length band structures shown in Fig. 2, indicating that BN coupling primarily affects boundary-localized states in the same-topology case.
Figure 3(b) shows four in-gap energy levels originating from two ESs and two IFs in the 9-7-9 AGNRH segment. The IFs arise from the effective end states of the central 7-AGNR segment with $R_{2}$ unit cells, as discussed in our previous work [DavidKuo, 12]. Within the physically relevant range $0.8~t\leq t_{in}\leq t$, corresponding to realistic BN coupling strengths, the IF levels $\Sigma_{IF,C(V)}$ are no longer well isolated from the bulk states. This behavior contrasts sharply with the $t_{in}=0$ limit and reflects the enhanced sublattice polarization induced by the symmetric BN environment, which drives the IF levels toward the bulk continuum.
The situation changes dramatically in the inverse-topology configuration shown in Figs. 3(c) and 3(d), where finite 7-AGNR and $9_{6}$-$7_{8}$-$9_{6}$ AGNRH segments are embedded between a BNNR and an NBNR. Although the bulk band gap ($E_{g}=E_{C}-E_{V}$) is substantially reduced for $0.8~t\leq t_{in}\leq t$, the energy levels associated with the ESs and IFs in the 9-7-9 AGNRH remain nearly unchanged and well separated from the bulk continuum. Such robustness is highly nontrivial and indicates that the inverse-topology configuration effectively preserves the IFs against strong environmental perturbations.
@@
To clarify the contrasting behaviors observed in Figs. 3(b) and 3(d), we plot the charge densities of $\Sigma_{IF,C}$ for different $t_{in}$ values in Fig. 4. Figures 4(a)-4(c) and Figs. 4(d)-4(f) correspond to the inverse-topology and same-topology configurations, respectively. For $t_{in}=0$, shown in Fig. 4(a), the AGNRH is completely decoupled from the BN environment, corresponding to vacuum boundary conditions. The state $\Sigma_{IF,C}$ originates from the coupling between a 9-7 junction IF state with A-sublattice chirality ($\Psi_{IF,A}(t_{in}=0)$) and a 7-9 junction IF state with B-sublattice chirality ($\Psi_{IF,B}(t_{in}=0)$). When a weak coupling $t_{in}=0.2~t$ is introduced, the charge density of $\Sigma_{IF,C}$ partially penetrates into the BN regions [Fig. 4(b)] while preserving mirror symmetry across the ribbon. Even for strong coupling $t_{in}=0.8~t$, as shown in Fig. 4(c), $\Sigma_{IF,C}$ maintains nearly equal contributions from $\psi_{IF,A}(t_{in}\neq 0)$ and $\psi_{IF,B}(t_{in}\neq 0)$. The charge-density distribution remains mirror-inverted across the ribbon due to the global inversion symmetry of the system. The increased energy separation between $\Sigma_{IF,C}$ and $\Sigma_{IF,V}$ is mainly attributed to the enhanced overlap between the wave functions $\psi_{IF,A}$ and $\psi_{IF,B}$. For finite $t_{in}$, the decay lengths of $\psi_{IF,A}$ and $\psi_{IF,B}$ increase as a result of interfacial-induced delocalization. Notably, charge densities simultaneously appear on both A- and B-sublattice sites at rows $i=5$ and $i=13$, reflecting the symmetric hybridization between the two junction states.
In contrast, Figs. 4(d)-4(f) demonstrate that in the same-topology configuration the IF charge densities become strongly polarized even at $t_{in}=0.2~t$, with dominant contributions from $\psi_{IF,A}$. The A(B)-sublattice IF states couple to nitride (boron) atoms at the lower and upper BN edges, respectively, leading to the formation of new bonding and antibonding states [PrunedaJM, 28]. As a result, $\Sigma_{IF,C(V)}$ shifts above (below) the mid-gap energy. This behavior indicates that the symmetric BN environment generates an effective field along the armchair direction, thereby breaking the inversion symmetry between the 9-7 and 7-9 junctions.
To further elucidate the influence of BN sheets on the IFs shown in Figs. 3 and 4, we introduce an effective Hamiltonian for the IF subspace:
$H_{eff,IF}$ (3)
$=$ $\left[\begin{array}[]{cc}E_{L}+\Sigma_{A}&t_{eff,LR}\\ t_{eff,LR}&E_{R}+\Sigma_{B}\\ \end{array}\right],$ (6)
where $E_{L}=0$ and $E_{R}=0$ denote the energy levels of the 9-7 and 7-9 junction IF states in the absence of $t_{in}$. $\Sigma_{A}$ and $\Sigma_{B}$ represent the self-energies induced by the interfacial hopping $t_{in}$.
For the inverse-topology configuration, the effective Hamiltonian reduces to
$H_{eff,IF,I}$ (7)
$=$ $\left[\begin{array}[]{cc}0&t_{eff,LR}\\ t_{eff,LR}&0\\ \end{array}\right],$ (10)
since the self-energies satisfy $\Sigma_{A(B)}=0$. More specifically, $\Sigma_{A(B)}=t_{in}^{2}(G_{Bo}+G_{Ni})$, where $G_{Bo}$ and $G_{Ni}$ are the Green’s functions describing electron propagation on boron and nitride atoms, respectively. Under the approximation $\Delta_{B}=-\Delta_{N}$, one has $G_{Bo}\approx-G_{Ni}$, leading to nearly vanishing self-energies. In the inverse-topology configuration, the upper BN ribbon is flipped relative to the lower one, reversing the BN sublattice alignment at the two interfaces and restoring chiral symmetry in the IF subspace.
For the same-topology configuration, the effective Hamiltonian becomes
$H_{eff,IF,S}$ (11)
$=$ $\left[\begin{array}[]{cc}2\Sigma_{Ni}&t_{eff,LR}\\ t_{eff,LR}&2\Sigma_{Bo}\\ \end{array}\right].$ (14)
In this case, the self-energies satisfy $\Sigma_{A(B)}=2\Sigma_{Ni(Bo)}$, which significantly breaks the chiral symmetry of the IF states. For the same-topology configuration, A sites in graphene couple to the same BN species (nitride atoms) on both sides, while B sites couple to boron atoms on both sides. The resulting asymmetric on-site renormalization explains the polarization of the IF charge densities observed in Fig. 4. The eigenstates of $H_{eff,IF,I}$ and $H_{eff,IF,S}$ therefore provide a consistent interpretation of the numerical charge-density distributions.
@@
The transmission spectra in the vicinity of the IF resonances are well described by ${\cal T}_{TDQD}(\varepsilon)=\frac{4\Gamma_{e,L}t^{2}_{eff,LR}\Gamma_{e,R}}{|(\varepsilon-E_{L}+i\Gamma_{e,L})(\varepsilon-E_{R}+i\Gamma_{e,R})-t^{2}_{eff,LR}|^{2}}$ in the absence of inter-dot and intra-dot Coulomb interactions [DavidKuo2, 38]. Here, $\Gamma_{e,L(R)}$ denotes the effective tunneling rate between the electrodes and the left (right) TQD, $E_{L(R)}=0$ represent the on-site energies of the two IF states, and $t_{eff,LR}$ is the effective interdot hopping strength. The zero-temperature electrical conductance is given by $G_{e}(\mu)=(2e^{2}/h){\cal T}_{TDQD}(\varepsilon=\mu)$, which reaches the conductance quantum $G_{0}=2e^{2}/h$ when $\mu$ aligns with the energy levels $\Sigma_{IF,C}$ and $\Sigma_{IF,V}$. Here, $\mu$ denotes the Fermi energy of the electrodes.
Notably, the spectra of $G_{e}(\mu)$ indicate that the effective hopping $t_{eff,LR}$ between the IFs in inverse-topology BN environments is enhanced compared with the vacuum boundary condition. This enhancement underscores the constructive role of topology-engineered environments in stabilizing the IFs and strengthening coherent topological transport. In Appendix A, we further investigate the electronic properties of a 15-13-15-AGNRH segment embedded in BN sheets, demonstrating that the above conclusions are not restricted to the 9-7-9 geometry.